OpenStudy (lgbasallote):

they say \[\sqrt[4]{16} = \pm 2\] however... \[\sqrt[4]{16(1)} \implies \sqrt[4]{16i^4} \implies \pm 2i\] where is the fallacy?

OpenStudy (anonymous):

no they do not

OpenStudy (anonymous):

the number 16 has 4 "fourth roots" namely solutions to \(x^4=16\) but only one that is represented as \(\sqrt[4]{16}\) which means the positive real one, namely 2

OpenStudy (lgbasallote):

so the others are -2, +2i, -2i?

OpenStudy (unklerhaukus):

oh there are four solutions

OpenStudy (lgbasallote):

yes...that makes sense...

OpenStudy (anonymous):

we consider \(y=\sqrt[4]{x}\) as a function for any given \(x\) u have just one \(y\) so \(\sqrt[4]{16}=2\)

OpenStudy (lgbasallote):

how would you explain what i wrote then @mukushla

OpenStudy (anonymous):

there is another error in reasoning here so say \(\sqrt[4]{ab}=\sqrt[4]{a}\sqrt[4]{b}\) you are assuming that \(a, b\) are positive real numbers

OpenStudy (lgbasallote):

so they're not?

OpenStudy (anonymous):

otherwise it is not true

OpenStudy (anonymous):

if you are living in the world of positive numbers, then yes and don't forget the symbol \(\sqrt{a}\) does not mean "some number whose square is \(a\)" but rather is means "the positive number whose square is \(a\)"

OpenStudy (lgbasallote):

i guess i really do have an engineer's mind and not a mathematician's =))

OpenStudy (unklerhaukus):